Write a formula for the an in each of the following sequences {an} n=1 given by

(i) 2, 1, 4, 3, 6, 5, 8, 7, . . .

(ii) 1, 3, 6, 10, 15, . . .

(iii) 1, −4, 9, −16, 25, −36, . . . Which ones among these are the subsequence of {n} n=1?


1
Expert's answer
2021-01-21T17:17:00-0500

Formulas for a given sequences:

1) "a_n=a_{n-1}+1+2(-1)^{n-1}" , "a_{1}=2"

and is not a subsequence of "{n}_{n=1}^{\infty}" cause have different order of members

2) "a_n=a_{n-1}+n" , "a_{1}=1"

and is a subsequence of "{n}_{n=1}^{\infty}"

3) "a_n=(-1)^{n-1}n^2"

and is not subsequence of "{n}_{n=1}^{\infty}" cause have other members and order of them